# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ A.2 Extension to Continuous Settings when $\mathcal{X}$ is Finite and $\Omega$ is Continuous
Assume that Regularity assumption, Assumption A.3, A.4 and A.5, and (A.1) hold. Let $\tau^{\dagger}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil$, and let $\Omega_{t}$ be a set of discretization for $\Omega$ with each coordinate equally divided into $\tau^{\dagger}_{t}$. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(2)}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil^{d_{2}}|\mathcal{X}|$, and define $\beta_{t}=2\log\kappa^{(2)}_{t}+\xi_{t}$. Let $q_{2}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.4 and A.5, respectively. Then, the following holds if Algorithm 3 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
@@
Under the assumption in Theorem A.3, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
@@
Under the assumption in Theorem A.3, the following holds:
@@
where $\hat{t}$ is given by (A.4), and functions $q_{2}(\cdot)$, $h^{\dagger}_{i}(\cdot)$ and each coefficient are given in Theorem A.3. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
Proofs are given by using the same argument as in the proof of Theorem 4.2 and 4.3. Finally, we consider $\hat{t}$ in the expectation measure. Under the expectation measure, since $\hat{\bm{x}}^{\dagger}_{t}$ corresponds to the posterior mean of $F^{\dagger}_{t}({\bm{x}})$, there is a gap with the index $\hat{t}$ given in (A.4). As a result, even in the case of the expectation measure, $\hat{t}$ does not necessarily equal $t$. Nevertheless, the upper bound of $\mathbb{E}[r^{\dagger}_{t}]$ can be expressed as the right-hand side in (A.5) plus $2t^{-2}$.
@@
Under the assumption in Theorem A.3, for the expectation measure, the following holds:
@@ A.3 Extension to Continuous Settings when $\mathcal{X}$ and $\Omega$ are Continuous
Algorithm 4 RRGP-UCB for robustness measures when $\mathcal{X}$ and $\Omega$ are continuous
@@
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (A.3)
   Select the next evaluation point $\bm{w}_{t}$ by ${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}})$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
@@ A.3.1 Theoretical Analysis in the Continuous Setting when $\mathcal{X}$ and $\Omega$ are Continuous
In order to derive the theoretical guarantee, we introduce the following two assumptions.
@@
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f({\bm{x}},[{\bm{w}}]_{t}))|$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}},[{\bm{w}}]_{t})|\right),$
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f([{\bm{x}}]_{t},[{\bm{w}}]_{t}))|$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f([{\bm{x}}]_{t},[{\bm{w}}]_{t})|\right)$
@@
For Assumption A.7, we can use $q(a)$ in Table 4 as $q_{3}(a)$ if the target robustness measure is given in Table 4. Details are described in Appendix C.5. Then, the following theorem holds.
@@
Assume that Regularity assumption, Assumption A.5, A.6 and A.7, and (A.1) hold. Let $\tilde{\tau}_{t}=\lceil b_{3}drt^{2}(\sqrt{\log(a_{3}d)}+\sqrt{\pi}/2)\rceil$, and let $\mathcal{X}_{t}\times\Omega_{t}$ be a set of discretization for $\mathcal{X}\times\Omega$ with each coordinate equally divided into $\tilde{\tau}_{t}$. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(3)}_{t}=(1+\tilde{\tau}^{d_{1}}_{t})\tilde{\tau}^{d_{2}}_{t}$, and define $\beta_{t}=2\log\kappa^{(3)}_{t}+\xi_{t}$. Let $q_{3}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.7 and A.5, respectively. Then, the following holds if Algorithm 4 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
@@
Under the assumption in Theorem A.6, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
@@
Under the assumption in Theorem A.6, the following holds:
@@
where $\hat{t}$ is given by (A.4), and functions $q_{3}(\cdot)$, $h^{\dagger}_{i}(\cdot)$ and each coefficient are given in Theorem A.6. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
The proof is given by using the same argument as in the proof of Theorem 4.2 and 4.3. Finally, for the expectation measure, the following theorem holds.
@@
Under the assumption in Theorem A.6, for the expectation measure, the following holds:
@@ B.1 Extension to Uncontrollable Settings when $\Omega$ is Finite
Let $\Omega$ be a finite set. In this case, if $\mathcal{X}$ is finite or continuous, the only difference between the proposed method and Algorithm 1 or 2 is whether or not ${\bm{w}}_{t}$ is sampled from $P^{\ast}$. Moreover, if we set $\kappa^{(1)}_{t}=|\mathcal{X}\times\Omega|$ in Algorithm 2, we can express the case when $\mathcal{X}$ is finite, i.e., Algorithm 1. We give the pseudo-code for the case where $\Omega$ is finite in the uncontrollable setting in Algorithm 5. Note that the $\mathcal{X}$ in Algorithm 5 includes both the finite and continuous cases.
Algorithm 5 RRGP-UCB for robustness measures in the uncontrollable setting when $\Omega$ is finite
@@
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (3.4)
@@
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
@@ B.2 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Finite
Let $\Omega=\{{\bm{w}}^{(1)},\ldots,{\bm{w}}^{(J)}\}$, $p_{j}=\mathcal{P}_{{\bm{w}}}({\bm{w}}={\bm{w}}^{(j)})$ and $p_{min}=\min_{1\leq j\leq J}p_{j}$. Next, we introduce the following assumption.
@@
Assume that Regularity assumption, Assumption 4.1 and B.1, and (3.2) hold. Suppose that $\mathcal{X}$ and $\Omega$ are finite. Also suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t}$, $\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Let $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be a function satisfying Assumption 4.1. Then, under the uncontrollable setting, the following holds if Algorithm 5 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C^{\prime}_{1}=\frac{2p^{-1}_{min}}{\log(1+\sigma^{-2}_{{\rm noise}})}$ and $C_{2,\nu_{ij}}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following inequality holds:
@@
where $C$ is $8$ for the mean absolute deviation measure and $4$ for other measures.
@@
Under the assumption in Theorem B.1, the following holds:
@@
where $\hat{t}$ is given by (4.3), and the function $h_{i}(\cdot)$ and each coefficient are given in Theorem B.1. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
Assume that Regularity assumption, Assumption 4.1, A.1, A.2 and B.1, and (3.2) hold. Suppose that $\mathcal{X}$ and $\Omega$ are continuous and finite, respectively. Also suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(4)}_{t}=(1+\lceil b_{1}d_{1}rt^{2}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)\rceil^{d_{1}})|\Omega|$, and define $\beta_{t}=2\log\kappa^{(4)}_{t}+\xi_{t}$. Let $q_{1}(a)$ and $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.2 and 4.1, respectively. Then, under the uncontrollable setting, the following holds if Algorithm 5 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C^{\prime}_{1}=\frac{2p^{-1}_{min}}{\log(1+\sigma^{-2}_{{\rm noise}})}$ and $C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.

### v2

@@ A.2 Extension to Continuous Settings when $\mathcal{X}$ is Finite and $\Omega$ is Continuous
Assume that the regularity assumption, Assumptions A.3, A.4 and A.5, and (A.1) hold. Let $\tau^{\dagger}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil$, and let $\Omega_{t}$ be a set of discretization for $\Omega$ with each coordinate equally divided into $\tau^{\dagger}_{t}$. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(2)}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil^{d_{2}}|\mathcal{X}|$, and define $\beta_{t}=2\log\kappa^{(2)}_{t}+\xi_{t}$. Let $q_{2}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumptions A.4 and A.5, respectively. Then, if Algorithm 4 is performed, the following holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
@@
Under the assumptions of Theorem A.3, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for the other measures.
@@
Under the assumptions of Theorem A.3, the following holds:
@@
where $\hat{t}$ is given by (A.4), and functions $q_{2}(\cdot)$, $h^{\dagger}_{i}(\cdot)$ and all coefficients are as defined in Theorem A.3. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
Proofs are given by using the same argument as in the proof of Theorems 4.2 and 4.3. Finally, we consider $\hat{t}$ in the expectation measure. Under the expectation measure, since $\hat{\bm{x}}^{\dagger}_{t}$ corresponds to the posterior mean of $F^{\dagger}_{t}({\bm{x}})$, there is a gap with the index $\hat{t}$ given in (A.4). As a result, even in the case of the expectation measure, $\hat{t}$ does not necessarily equal $t$. Nevertheless, the upper bound of $\mathbb{E}[r^{\dagger}_{t}]$ can be expressed as the right-hand side in (A.5) plus $2t^{-2}$.
@@
Under the assumptions of Theorem A.3, for the expectation measure, the following holds:
@@ A.3 Extension to Continuous Settings when $\mathcal{X}$ and $\Omega$ are Continuous
Algorithm 5 RRGP-UCB for robustness measures when $\mathcal{X}$ and $\Omega$ are continuous.
@@
   Generate $\xi_{t}$ from chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu^{\dagger}_{t-1}({\bm{x}},\cdot)})$
   Select next evaluation point $\bm{x}_{t}$ by (A.3)
   Select next evaluation point $\bm{w}_{t}$ by ${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}})$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding observed data
@@ A.3.1 Theoretical Analysis in the Continuous Setting when $\mathcal{X}$ and $\Omega$ are Continuous
To derive the theoretical guarantee, we introduce the following two assumptions.
@@
$|\rho({f({\bm{x}},\cdot)})-\rho({f^{\dagger}_{t}({\bm{x}},\cdot)})|$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}},[{\bm{w}}]_{t})|\right),$
$|\rho({f({\bm{x}},\cdot)})-\rho({f^{\dagger}_{t}([{\bm{x}}]_{t},\cdot)})|$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f([{\bm{x}}]_{t},[{\bm{w}}]_{t})|\right)$
@@
For Assumption A.7, we can use $q^{(m)}(a)$ in Table 4 in Inatsu et al., 2024a as $q_{3}(a)$ if the target robustness measure is the Bayes risk (expectation), worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, or mean absolute deviation measures described in the table. Details are described in Appendix C.5. Then, the following theorem holds.
@@
Assume that the regularity assumption, Assumptions A.5, A.6 and A.7, and (A.1) hold. Let $\tilde{\tau}_{t}=\lceil b_{3}drt^{2}(\sqrt{\log(a_{3}d)}+\sqrt{\pi}/2)\rceil$, and let $\mathcal{X}_{t}\times\Omega_{t}$ be a set of discretization for $\mathcal{X}\times\Omega$ with each coordinate equally divided into $\tilde{\tau}_{t}$. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(3)}_{t}=(1+\tilde{\tau}^{d_{1}}_{t})\tilde{\tau}^{d_{2}}_{t}$, and define $\beta_{t}=2\log\kappa^{(3)}_{t}+\xi_{t}$. Let $q_{3}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumptions A.7 and A.5, respectively. Then, if Algorithm 5 is performed, the following holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
@@
Under the assumptions of Theorem A.6, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for the other measures.
@@
Under the assumptions of Theorem A.6, the following holds:
@@
where $\hat{t}$ is given by (A.4), and functions $q_{3}(\cdot)$, $h^{\dagger}_{i}(\cdot)$ and all coefficients are as defined in Theorem A.6. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
The proof is given by using the same argument as in the proof of Theorems 4.2 and 4.3. Finally, for the expectation measure, the following theorem holds.
@@
Under the assumptions of Theorem A.6, for the expectation measure, the following holds:
@@ B.1 Extension to Uncontrollable Settings when $\Omega$ is Finite
Let $\Omega$ be a finite set. In this case, if $\mathcal{X}$ is finite or continuous, the only difference between the proposed method and Algorithm 1 or 3 is whether or not ${\bm{w}}_{t}$ is sampled from $P^{\ast}$. Moreover, if we set $\kappa^{(1)}_{t}=|\mathcal{X}\times\Omega|$ in Algorithm 3, we can express the case when $\mathcal{X}$ is finite, i.e., Algorithm 1. We give the pseudo-code for the case where $\Omega$ is finite in the uncontrollable setting in Algorithm 6. Note that the $\mathcal{X}$ in Algorithm 6 includes both the finite and continuous cases.
Algorithm 6 RRGP-UCB for robustness measures in the uncontrollable setting when $\Omega$ is finite.
@@
   Generate $\xi_{t}$ from chi-squared distribution with two degrees of freedom
@@
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu_{t-1}({\bm{x}},\cdot)})$
   Select next evaluation point $\bm{x}_{t}$ by (3.4)
@@
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding observed data
@@ B.2 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Finite
Let $\Omega=\{{\bm{w}}^{(1)},\ldots,{\bm{w}}^{(J)}\}$, $p_{j}=\mathbb{{P}}_{{\bm{w}}}({\bm{w}}={\bm{w}}^{(j)})$ and $p_{min}=\min_{1\leq j\leq J}p_{j}$. Next, we introduce the following assumption.
@@
Assume that the regularity assumption, Assumptions 4.1 and B.1, and (3.2) hold. Suppose that $\mathcal{X}$ and $\Omega$ are finite. Also suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t}$, $\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(4)}_{t}=|\mathcal{X}\times\Omega|$, and define $\beta_{t}=2\log\kappa^{(4)}_{t}+\xi_{t}$. Let $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be a function satisfying Assumption 4.1. Then, under the uncontrollable setting, if Algorithm 6 is performed, the following holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C^{\prime}_{1}=\frac{2p^{-1}_{min}}{\log(1+\sigma^{-2}_{{\rm noise}})}$, $C_{2,\nu_{ij}}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following inequality holds:
@@
where $C$ is $8$ for the mean absolute deviation measure and $4$ for the other measures.
@@
Under the assumptions of Theorem B.1, the following holds:
@@
where $\hat{t}$ is given by (4.4), and the function $h_{i}(\cdot)$ and all coefficients are as defined in Theorem B.1. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
Assume that the regularity assumption, Assumptions 4.1, A.1, A.2 and B.1, and (3.2) hold. Suppose that $\mathcal{X}$ and $\Omega$ are continuous and finite, respectively. Also suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(4)}_{t}=(1+\lceil b_{1}d_{1}rt^{2}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)\rceil^{d_{1}})|\Omega|$, and define $\beta_{t}=2\log\kappa^{(4)}_{t}+\xi_{t}$. Let $q_{1}(a)$ and $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumptions A.2 and 4.1, respectively. Then, under the uncontrollable setting, if Algorithm 6 is performed, the following holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C^{\prime}_{1}=\frac{2p^{-1}_{min}}{\log(1+\sigma^{-2}_{{\rm noise}})}$, $C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following holds:
@@
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for the other measures.